OpenStudy (anonymous):

What is the maximum distance we can shoot a dart, provided our toy dart gun gives a maximum initial velocity of 5.80 m/s? the answer is 1.16m I just don't know why?!!!? help please!

OpenStudy (anonymous):

The formula for the distance is v^2*sin(2alfa)/2g. To get your object as far as you can, you need to shoot it at angular of 45 degree (because of sin(2*45)=1). So, D=(5.8^2*1)/(2*10)=1.628 meters.